OpenStudy (anonymous):

Solve for X: 12/x^2-16 - 24/x-4 = 3

OpenStudy (anonymous):

Using difference of squares:\[ (a^2-b^2) = (a-b)(a+b) \]So: \[ (x^2-16) = (x^2-4^2) = (x-4)(x+4) \]

OpenStudy (anonymous):

Multiplying both sides by \((x-4)(x+4)\) you'd get: \[ 12 - 24(x+4) = 3(x-4)(x+4) \]So then it's just a matter of simplifying and stuff.

OpenStudy (anonymous):

If you can get it in the form of \[ ax^2+bx+c = 0 \]Then your answer is just the quadratic formula: \[ x =\frac{-b \pm \sqrt{b^2-4ac}}{2a} \]